# Scaling Gaussian Process Regression with Full Derivative Observations

### v1

@@ A.2 Solving
$\alpha\leftarrow\texttt{solve}(\hat{\tilde{\mathbf{C}}}\alpha=\hat{\tilde{\mathbf{K}}}_{\mathbf{z}\mathbf{x}}\tilde{\mathbf{\Lambda}}^{-1}\tilde{\mathbf{y}})$ (34)
@@ A.3 Implementation
Recall that we use CG to implement the Hutchinson pseudoloss. We use a PyTorch implementation of CG descent method following (Maddox et al., 2022). We use a CG tolerance of $51e-$ 5. We use a rank-10 pivoted Cholseky decomposition to obtain a preconditioner for CG descent.
@@ Appendix B Additional Experimental Results
B.1 Effect of Noise Levels
The DSoftKI method balances the fitting of value information with gradient information by controlling the relative ratio of $\beta^{2}_{v}$ to $\beta^{2}_{g}$. Figure 5 shows the results of varying this ratio on the synthetic benchmark to gain more insight into how to set this ratio. We see that the relatative ratio of $\beta^{2}_{v}$ to $\beta^{2}_{g}$ is dependent on the dataset. Nevertheless, the general pattern is that increasing the weighting of the values tends to improve performance. The experiments reported in the paper had a relative factor of $1$.
@@ B.1 Effect of Noise Levels
(b) Six Hump Camel.
(c) Styblinski Tang.
@@
B.2 More Synthetic Benchmark Results
SoftKI SVGP DSVGP DDSVGP DSoftKI
Branin 2 0.171 $\pm$ 0.023 0.239 $\pm$ 0.035 0.378 $\pm$ 0.11 1.557 $\pm$ 0.078 0.075 $\pm$ 0.004
Six-hump-camel 2 1.067 $\pm$ 0.156 1.221 $\pm$ 0.055 2.465 $\pm$ 0.87 7.115 $\pm$ 0.211 0.348 $\pm$ 0.096
Styblinski-tang 2 1.008 $\pm$ 0.106 0.961 $\pm$ 0.04 0.871 $\pm$ 0.033 0.808 $\pm$ 0.22 0.322 $\pm$ 0.063
Hartmann 6 0.429 $\pm$ 0.232 0.377 $\pm$ 0.006 0.507 $\pm$ 0.011 0.527 $\pm$ 0.019 0.028 $\pm$ 0.008
Welch 20 0.011 $\pm$ 0.001 0.053 $\pm$ 0.002 - 0.269 $\pm$ 0.003 0.001 $\pm$ 0.001
Table 5: Gradient fitting errors on synthetic functions using scaled effective learning rate.
@@ B.2 More Synthetic Benchmark Results
Branin 0.037 $\pm$ 0.399 127.566 $\pm$ 48.658 0.393 $\pm$ 0.162
Six-hump-camel 0.415 $\pm$ 0.123 56.261 $\pm$ 53.971 7.946 $\pm$ 6.107
Styblinski-tang 0.651 $\pm$ 0.049 42.629 $\pm$ 15.41 5.335 $\pm$ 1.945
Hartmann 0.049 $\pm$ 0.013 10.135 $\pm$ 1.544 0.804 $\pm$ 0.171
Welch - -0.723 $\pm$ 2.583 0.228 $\pm$ 0.283
@@
We also compare the gradient test RMSE of the base GP and its extension with derivatives. For a gradient test RMSE, we sum up the errors across the dimensions. Thus, we should normalize by the numbers to compare across datasets. We observe that it is dataset dependent whether or not DSVGP outperforms SVGP. As a reminder, DDSVGP does not fit all the derivative information, and thus, we derive its gradient prediction from its value prediction. We observe that derivative information is helpful for DSoftKI, the extension of SoftKI to the setting with derivatives.
@@
(g) Double-Walled Nanotube.
@@
(g) Double-Walled Nanotube.
@@
B.3 More Force Field Results
@@
Training curves. Figure 7 and Figure 8 give the test RMSE, test gradient RMSE, and their standard deviations at each epoch of training on the MD22 dataset. As a reminder, we did not tune the learning rates or batch sizes to obtain the best possible performance. Rather, we control the effective learning rate so that we can compare GP and GPwD regression. SVGP is in blue, SoftKI is in green, DDSVGP is in orange, and DSoftKI is in Red. On Ac-Ala3-NHMe, we observe that methods that fit derivative information overfit both values and gradients. On several datasets such as AT-AT and Stachyose, we observe that DSoftKI overfits on test RMSE but does not on test gradient RMSE. On other dataset such as AT-AT-CG-CG, Buckyball-catcher, and Double-walled nanotube, we observe that DSoftKI is able to fit both the values and gradients. In contrast, DDSVGP tends to overfit as measured by test gradient RMSE. This is somewhat surprising, since it does fit directional derivative information. We also observe that SoftKI and SVGP cannot fit gradients even though they do fit the values. This suggests that GPs that do not fit derivatives do not capture the shape of the surface well.
Lengthscales and temperatures. Figure 9 and Figure 10 give the histograms of lengthscales and temperatures that are learned by DSoftKI and SoftKI. We see that the lengthscales that are learned by DSoftKI and SoftKI are different, giving further evidence that interpolation points in the derivative and non-derivative versions are utilized differently. In particular, the lengthscales learned by the DSoftKI tend to be larger than the lengthscales learned by SoftKI. As a reminder, the kernel $\tilde{\mathbf{K}}_{\mathbf{z}\mathbf{z}}(\ell)$ is a function of the lengthscales $\ell$ and not the temperature. However, the temperatures that are learned are similar. Similarly, the interpolation $\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}(T)$ is a function of the temperatures and not the lengthscales. This suggests that the interpolation is utilized similarly across the derivative and non-derivative versions.
@@
(g) Double-Walled Nanotube.
@@
(g) Double-Walled Nanotube.
Appendix C Background on DDSVGP and DSKI

### v2

@@ Appendix A Method
We describe the DSoftKI method in more detail, including comparing its kernel structure with related GPwD methods (Section A.1), detailing the solving procedure (Section A.2), and its implementation (Section A.3).
@@ A.2 Solving
[…]
@@ A.3 Implementation
Recall that we use CG to implement the Hutchinson pseudoloss. We use a PyTorch implementation of CG descent method following (Maddox et al., 2022). We use a CG tolerance of 1e51e-5. We use a rank-10 pivoted Cholseky decomposition to obtain a preconditioner for CG descent.
Test RMSE Test NLL
d Exact MLL Hutchinson Exact MLL Hutchinson
Branin 2 0.002 $\pm$ nan 0.029 $\pm$ 0.0 -4.527 $\pm$ nan 0.208 $\pm$ 0.003
Six-hump-camel 2 0.014 $\pm$ 0.002 0.198 $\pm$ 0.018 -2.185 $\pm$ 0.245 0.627 $\pm$ 0.014
Styblinski-tang 2 0.013 $\pm$ 0.004 0.035 $\pm$ 0.001 -2.283 $\pm$ 0.065 0.726 $\pm$ 0.08
Hartmann 6 0.011 $\pm$ 0.003 0.408 $\pm$ 0.016 -3.145 $\pm$ 0.318 1.159 $\pm$ 0.007
Welch 20 - 0.618 $\pm$ 0.006 - 1.373 $\pm$ 0.002
Table 7: Comparison of using Exact MLL vs. Hutchinson’s pseudoloss. The best test RMSE and test NLL are bolded. The notation $\pm\texttt{nan}$ indicates that one run was unstable whereas the notation - indicates that all runs failed.
@@ Appendix B Additional Experimental Results
We detail the hyperparameters used (Section B.1). We also provide supplemental ablations (Section B.2 and Section B.3) and supplemental experimental results for the experiments described in the main text (Section B.4 and Section B.5).
B.1 Hyperparameters
For the experiments in this paper, we initialize hyperparameters as follows:
  • lengthscales $\ell_{j}=1$,
  • output scale $\gamma=1$,
  • interpolation points $\mathbf{z}$ via $k$-means clustering on training data,
  • temperatures $T_{kj}=1$ for all entries,
  • value noise $\beta^{2}_{v}=0.1$, and
  • gradient noise $\beta^{2}_{g}=0.1\times d$ where $d$ is the dimensionality of the dataset.
As discussed in the main text, all hyperparameters
$\theta=(\ell,\gamma,\mathbf{z}^{(m\times d)},\mathbf{T}^{(m\times d)},\beta_{v},\beta_{g})$ (40)
are learned jointly via gradient-based optimization of the stabilized DSoftKI MLL (Equation 25) using Adam (Kingma and Ba, 2014).
B.2 Exact MLL vs. Hutchinson’s Pseudoloss.
Table 7 reproduces the results of Section 5.1 using the Exact MLL compared to Hutchinson’s pseudoloss. We see that when using DSoftKI’s MLL, there is instability in training on Branin and Welch. However, using Hutchionson’s pseudoloss exclusively leads to worse results. As a result, we opt for an objective that uses the DSoftKI MLL, when possible, and Hutchinson’s pseudoloss otherwise as described in Section 4.2.
B.3 Effect of Noise Levels
The DSoftKI method balances the fitting of value information with gradient information by controlling the relative ratio of $\beta^{2}_{v}$ to $\beta^{2}_{g}$. Figure 5 shows the results of varying this ratio on the synthetic benchmark to gain more insight into how to set this ratio. We see that the relatative ratio of $\beta^{2}_{v}$ to $\beta^{2}_{g}$ is dependent on the dataset. The experiments reported in the paper had a relative factor of $1$.
@@ B.3 Effect of Noise Levels
(b) Six-hump-camel.
(c) Styblinski-tang.
@@
B.4 More Synthetic Benchmark Results
d SoftKI SVGP DSVGP DDSVGP DSoftKI
Branin 2 0.17 $\pm$ 0.024 0.237 $\pm$ 0.029 0.357 $\pm$ 0.207 1.526 $\pm$ 0.058 0.07 $\pm$ 0.01
Six-hump-camel 2 1.069 $\pm$ 0.156 1.206 $\pm$ 0.057 1.545 $\pm$ 0.071 7.032 $\pm$ 0.187 0.345 $\pm$ 0.084
Styblinski-tang 2 1.008 $\pm$ 0.11 0.998 $\pm$ 0.02 0.86 $\pm$ 0.133 0.895 $\pm$ 0.165 0.319 $\pm$ 0.057
Hartmann 6 0.429 $\pm$ 0.232 0.377 $\pm$ 0.007 0.506 $\pm$ 0.01 0.527 $\pm$ 0.01 0.028 $\pm$ 0.008
Welch 20 0.011 $\pm$ nan 0.052 $\pm$ 0.0 - 0.269 $\pm$ 0.002 0.001 $\pm$ 0.001
Table 8: Gradient fitting errors on synthetic functions using scaled effective learning rate. One of the runs for SoftKI encountered numerical instability (hence $\pm\texttt{nan}$). - indicates a timeout for DSVGP.
@@ B.4 More Synthetic Benchmark Results
Branin 0.323 $\pm$ 0.821 281.505 $\pm$ 72.881 0.397 $\pm$ 0.061
Six-hump-camel 0.33 $\pm$ 0.248 53.827 $\pm$ 49.856 8.921 $\pm$ 7.412
Styblinski-tang 0.665 $\pm$ 0.121 41.907 $\pm$ 14.305 4.929 $\pm$ 2.126
Hartmann 0.047 $\pm$ 0.014 10.755 $\pm$ 0.916 0.591 $\pm$ 0.265
Welch - -1.329 $\pm$ 1.906 0.138 $\pm$ 0.161
@@
We also compare the gradient test RMSE of the base GP and its extension with derivatives. For a gradient test RMSE, we sum up the errors across the dimensions. Thus, we should normalize by the number of dimensions to compare across datasets. We observe that it is dataset dependent whether or not DSVGP outperforms SVGP. As a reminder, DDSVGP does not fit all the derivative information, and thus, we derive its gradient prediction from its value prediction. We observe that derivative information is helpful for DSoftKI, the extension of SoftKI to the setting with derivatives.
@@
(g) Double-walled-nanotube.
@@
(g) Double-walled-nanotube.
@@
B.5 More Force Field Results
@@
Training curves. Figure 7 and Figure 8 give the test RMSE, test gradient RMSE, and their standard deviations at each epoch of training on the MD22 dataset. As a reminder, we did not tune the learning rates or batch sizes to obtain the best possible performance. Rather, we control the effective learning rate so that we can compare GP and GPwD regression. SVGP is in blue, SoftKI is in green, DDSVGP is in orange, and DSoftKI is in Red. On Ac-Ala3-NHMe, we observe that methods that fit derivative information overfit both values and gradients. On several datasets such as AT-AT and Stachyose, we observe that DSoftKI overfits on test RMSE but does not on test gradient RMSE. On other dataset such as AT-AT-CG-CG, Buckyball-catcher, and Double-walled-nanotube, we observe that DSoftKI is able to fit both the values and gradients. In contrast, DDSVGP tends to overfit as measured by test gradient RMSE. This is somewhat surprising, since it does fit directional derivative information. We also observe that SoftKI and SVGP cannot fit gradients even though they do fit the values. This suggests that GPs that do not fit derivatives do not capture the shape of the surface well.
Temperatures and lengthscales. Figure 9 and Figure 10 give the histograms of temperatures and lengthscales that are learned by DSoftKI and SoftKI. As a reminder, we modified SoftKI to also have different temperatures per interpolation point as we did for DSoftKI to have a more fair comparison.
The temperature $T_{k}\in\mathbb{R}^{d}$ for interpolation point $z_{k}$ controls how strongly that point responds to variations along each input dimension during interpolation. A smaller temperature component $T_{kj}$ makes the interpolation weight $\sigma_{z}^{j}(x)$ more sensitive to changes in dimension $j$, effectively sharpening the directional influence. Conversely, larger temperatures smooth out the response along that dimension.
Figure 9 shows that the learned temperature distributions are similar between DSoftKI and SoftKI, indicating that the interpolation scheme (which is what is affected by the temperature) is primarily adapted to the data geometry. Moreover, we observe that DSoftKI learns larger kernel lengthscales compared to SoftKI (Figure 10). This suggests that when gradient information is available, the model can rely on the temperature vectors to capture local directional variation, allowing the kernel lengthscales to increase and model smoother global structure. In other words, the temperatures and lengthscales play complementary roles: temperatures handle local directional sensitivity while lengthscales control global smoothness.
@@
(g) Double-walled-nanotube.
@@
(g) Double-walled-nanotube.
Appendix C Additional Datasets
We describe experiments on additional datasets including a toy n-body simulation (Section C.1) and the UCI dataset with synthetic gradients (Section C.2) to test DSoftKI on a broader range of datasets.
C.1 Toy N-Body Dataset
d DDSVGP DSoftKI $\nabla$-DDSVGP $\nabla$-DSoftKI
nbody-4 24 0.058 $\pm$ 0.005 0.059 $\pm$ 0.002 1.453 $\pm$ 0.083 0.708 $\pm$ 0.034
nbody-6 36 0.027 $\pm$ 0.001 0.035 $\pm$ 0.002 1.5 $\pm$ 0.059 1.032 $\pm$ 0.045
nbody-8 48 0.03 $\pm$ 0.0 0.028 $\pm$ 0.002 1.368 $\pm$ 0.117 1.016 $\pm$ 0.092
nbody-10 60 0.026 $\pm$ 0.003 0.024 $\pm$ 0.0 1.343 $\pm$ 0.037 1.104 $\pm$ 0.04
Table 10: Test RMSE (best bolded) on $n$-body datasets. $\nabla$-DDSVGP and $\nabla$-DSoftKI denote gradient RMSE (best bolded).
d DDSVGP DSoftKI
nbody-4 24 0.98 $\pm$ 0.016 1.68 $\pm$ 0.019
nbody-6 36 0.995 $\pm$ 0.029 2.802 $\pm$ 0.048
nbody-8 48 0.97 $\pm$ 0.006 3.646 $\pm$ 0.059
nbody-10 60 0.976 $\pm$ 0.001 4.591 $\pm$ 0.162
Table 11: Training time per epoch (in seconds, best bolded) on $n$-body Hamiltonian datasets.
Computational physics and chemistry provide settings where high-dimensional gradient observations arise naturally. Consider an $n$-body system with $n$ particles (e.g., atoms or point masses). Let $\mathbf{q}=(q_{1},\dots,q_{n})$ denote the positions of each particle where each $q_{i}\in\mathbb{R}^{3}$, and let $\mathbf{p}=(p_{1},\dots,p_{n})$ denote the momenta where each $p_{i}\in\mathbb{R}^{3}$. The dynamics are governed by Hamilton’s equations
$\frac{d\mathbf{q}}{dt}=\frac{\partial H}{\partial\mathbf{p}},\quad\frac{d\mathbf{p}}{dt}=-\frac{\partial H}{\partial\mathbf{q}}$ (41)
where $H(\mathbf{q},\mathbf{p})$ is the Hamiltonian. Thus, given observations of the Hamiltonian $H$ and its gradients $\nabla H$, we can learn the system’s dynamics, i.e., $\frac{d\mathbf{q}}{dt}$ and $\frac{d\mathbf{p}}{dt}$. More concretely, it is a GPwD regression problem where $\mathbf{x}=(\mathbf{q},\mathbf{p})$, $\mathbf{y}=H(\mathbf{x})$, and
$\nabla\mathbf{y}$ $=\left(\frac{\partial H}{\partial\mathbf{q}},\frac{\partial H}{\partial\mathbf{p}}\right)\,.$ (42)
The input dimensionality is $d=6n$ since each particle contributes $3$ position and $3$ momentum coordinates. A learned surrogate for the Hamiltonian enables efficient prediction of the system’s dynamics without expensive numerical integration, which can be intractable for large $n$.
We construct an $n$-body gravitational dataset $\texttt{nbody-}n$ as follows. [^3] This is a classical system, in contrast to the MD22 dataset (Section 5.2) which is a quantum system. The Hamiltonian is
$H(\mathbf{q},\mathbf{p})=\sum_{i=1}^{n}\frac{\|p_{i}\|^{2}}{2m_{i}}-\sum_{i<j}\frac{Gm_{i}m_{j}}{\sqrt{\|q_{i}-q_{j}\|^{2}+\epsilon^{2}}}$ (43)
where $m_{i}$ is the mass of particle $i$, $G$ is the gravitational constant, and $\epsilon$ is a softening parameter that prevents singularities when particles approach each other. The first term is the kinetic energy and the second is the gravitational potential energy with Plummer softening (Plummer, 1911). In this toy example, the Hamiltonian has analytic gradients which may not exist for more complex systems.
We generate datasets for $n\in\{4,6,8,10\}$ particles, corresponding to input dimensions $d\in\{24,36,48,60\}$. To generate physically realistic configurations, we simulate $100$ trajectories by integrating Hamilton’s equations using a high-order Runge-Kutta method (DOP853) as implemented in SciPy (Virtanen et al., 2020). Initial positions are sampled from $\mathcal{N}(0,2^{2})$, initial momenta from $\mathcal{N}(0,0.5^{2})$ scaled by particle mass, and masses uniformly from $[0.5,2.0]$. We apply center-of-mass corrections for numerical stability. From each trajectory, we sample configurations to obtain $10000$ total samples per setting, computing the Hamiltonian and its gradients analytically. We filter out the 5% of samples with the largest gradient norms for numerical stability, leaving $9500$ samples. For this toy dataset, we use $G=1$ and $\epsilon=0.1$.
Table 10 reports the results of using a 90/10 training-testing split on each dataset. DSoftKI and DDSVGP achieve similar value RMSE, but DSoftKI substantially outperforms DDSVGP on gradient prediction (e.g., 0.708 RMSE versus 1.453 $\nabla$-RMSE for $n=4$). This demonstrates DSoftKI’s advantage in applications requiring accurate full gradient predictions, such as learning physical dynamics from energy observations, and justifies the higher computational cost (Table 11).
C.2 UCI Dataset
n d DDSVGP DSoftKI $\nabla$-DDSVGP $\nabla$-DSoftKI
Kin40k 18000 8 0.867 $\pm$ 0.002 0.864 $\pm$ 0.006 0.399 $\pm$ 0.001 0.393 $\pm$ 0.001
Protein 20578 9 0.81 $\pm$ 0.006 0.787 $\pm$ 0.003 2.778 $\pm$ 0.087 2.655 $\pm$ 0.283
Bike 7820 17 0.629 $\pm$ 0.026 0.421 $\pm$ 0.006 0.763 $\pm$ 0.004 0.624 $\pm$ 0.012
Elevators 7470 18 0.821 $\pm$ 0.013 0.797 $\pm$ 0.009 0.449 $\pm$ 0.002 0.448 $\pm$ 0.001
Pol 6750 26 0.898 $\pm$ 0.007 0.702 $\pm$ 0.01 0.383 $\pm$ 0.013 0.324 $\pm$ 0.012
Slice 19260 385 0.357 $\pm$ 0.005 0.03 $\pm$ 0.008 0.165 $\pm$ 0.007 0.094 $\pm$ 0.004
Table 12: Test RMSE (best bolded) on selected UCI datasets. $\nabla$-DDSVGP and $\nabla$-DSoftKI denote gradient RMSE (best bolded).
n d DDSVGP DSoftKI
Kin40k 18000 8 2.257 $\pm$ 0.022 1.715 $\pm$ 0.1
Protein 20578 9 2.593 $\pm$ 0.05 2.054 $\pm$ 0.067
Bike 7820 17 0.977 $\pm$ 0.003 1.317 $\pm$ 0.067
Elevators 7470 18 0.994 $\pm$ 0.024 1.297 $\pm$ 0.017
Pol 6750 26 0.876 $\pm$ 0.014 1.663 $\pm$ 0.077
Slice 19260 385 2.378 $\pm$ 0.047 71.279 $\pm$ 0.076
Table 13: Wall-clock training time (in seconds) per epoch (best bolded) on selected UCI datasets.
To evaluate scaling on diverse input distributions beyond physics datasets, we construct GPwD regression benchmarks from standard UCI regression datasets (Dua and Graff, 2017) by generating synthetic gradients. We select six UCI datasets spanning a range of dimensions: Kin40k ($d=8$), Protein ($d=9$), Bike ($d=17$), Elevators ($d=18$), Pol ($d=26$), and Slice ($d=385$). For each dataset, we estimate gradients using a $k$-nearest neighbor finite difference approximation: for each point $x_{i}$ with label $y_{i}$, we compute directional derivatives along the directions to its $k=3$ nearest neighbors and average the resulting gradient estimates, i.e.,
$\nabla y_{i}\approx\frac{1}{k}\sum_{j\in N_{k}(i)}\frac{y_{j}-y_{i}}{\|x_{j}-x_{i}\|^{2}}(x_{j}-x_{i})$ (44)
where $N_{k}(i)$ denotes the $k$ nearest neighbors of point $i$. This introduces noise into the gradient observations, providing a test of robustness to gradient estimation error.
Table 12 reports the results. DSoftKI outperforms DDSVGP on both value and gradient prediction across all datasets. The Slice dataset ($d=385$) is particularly notable: DSoftKI achieves 0.030 RMSE versus DDSVGP’s 0.357, demonstrating strong performance on high-dimensional structured data. The higher training time for DSoftKI on Slice (Table 13) reflects its $O(m^{2}nd)$ complexity, but is justified by substantially better accuracy.
Appendix D Deep Kernel Learning
DSoftKI DSoftKI-DKL
Ac-ala3-nhme 5.98e-03 $\pm$ 2.90e-05 6.93e-03 $\pm$ 9.40e-05
Dha 5.85e-03 $\pm$ 8.80e-05 6.19e-03 $\pm$ 7.40e-05
At-at 4.01e-03 $\pm$ 1.04e-04 5.63e-03 $\pm$ 5.40e-05
Stachyose 3.99e-03 $\pm$ 2.90e-05 5.29e-03 $\pm$ 1.30e-05
At-at-cg-cg 2.58e-03 $\pm$ 5.60e-05 3.98e-03 $\pm$ 2.14e-04
Buckyball-catcher 1.81e-03 $\pm$ 5.40e-05 4.09e-03 $\pm$ 1.17e-04
Double-walled-nanotube 9.33e-04 $\pm$ 2.70e-05 2.90e-03 $\pm$ 4.59e-04
Table 14: Normalized test RMSE per atom (best bolded) on MD22 dataset.
DSoftKI DSoftKI-DKL
Ac-ala3-nhme 8.27e-02 $\pm$ 1.73e-04 8.62e-02 $\pm$ 1.00e-05
Dha 7.41e-02 $\pm$ 2.59e-04 7.62e-02 $\pm$ 2.27e-04
At-at 7.03e-02 $\pm$ 8.50e-05 7.49e-02 $\pm$ 1.25e-04
Stachyose 5.86e-02 $\pm$ 6.00e-05 6.19e-02 $\pm$ 1.04e-04
At-at-cg-cg 5.08e-02 $\pm$ 2.02e-04 5.32e-02 $\pm$ 1.10e-04
Buckyball-catcher 4.09e-02 $\pm$ 1.78e-04 4.75e-02 $\pm$ 1.66e-04
Double-walled-nanotube 2.84e-02 $\pm$ 5.10e-05 3.28e-02 $\pm$ 1.04e-03
Table 15: Normalized gradient test RMSE per dimension $d$ (best bolded) on MD22 dataset.
n d DSoftKI DSoftKI-DKL
Ac-ala3-nhme 76598 126 82.665 $\pm$ 0.06 58.98 $\pm$ 0.516
Dha 62777 168 89.155 $\pm$ 0.223 48.736 $\pm$ 0.56
At-at 18000 180 27.236 $\pm$ 0.037 13.958 $\pm$ 0.025
Stachyose 24544 261 54.181 $\pm$ 0.142 19.324 $\pm$ 0.369
At-at-cg-cg 9137 354 27.102 $\pm$ 0.036 12.574 $\pm$ 0.339
Buckyball-catcher 5491 444 20.764 $\pm$ 0.111 7.655 $\pm$ 0.146
Double-walled-nanotube 4528 1110 45.379 $\pm$ 0.007 7.143 $\pm$ 0.088
Table 16: Comparing (wall-clock) training time (in seconds) per epoch (best bolded) on MD22 dataset for DSoftKI versus DSoftKI-DKL.
One advantage of DSoftKI over DSVGP and DDSVGP is that it does not require computing first or second-order derivatives of the kernel (Section 4.1). This enables the use of learned kernels, such as Deep Kernel Learning (Wilson et al., 2016) (DKL), without hard-coding kernel derivatives so that standard automatic differentiation techniques can be applied.
To demonstrate this capability, we evaluate DSoftKI with a deep kernel learning (DSoftKI-DKL) setup on the MD22 molecular datasets. We use a 2-layer MLP feature extractor with hidden dimension $64$, output dimension $24$, and tanh activations. The RBF kernel is applied in the learned feature space. We train all hyperparameters jointly using Adam, with a learning rate of $0.002$ for the GP hyperparameters and $0.0005$ for the neural network weights.
Table 14, Table 15, and Table 16 report the results. DSoftKI-DKL performs comparably to DSoftKI, with slightly higher RMSE on both values and gradients. We emphasize that this experiment is intended as a demonstration that deep kernel learning works out of the box with DSoftKI, rather than a claim that it improves performance on these particular datasets. Notably, DSoftKI-DKL offers computational speedups over DSoftKI since the interpolation points are projected into a lower-dimensional feature space. Furthermore, the neural network gradients can be computed efficiently via Jacobian-vector products, whereas DSVGP and DDSVGP would require hard-coding the feature extractor Jacobians or face intractable computation.
Appendix E Additional Background on DDSVGP and DSKI